OpenStudy (anonymous):

HELP!!! A circle has a diameter of 14 ft. If the area is tripled, what happens to the circumference?

OpenStudy (campbell_st):

the circumference will increase by a factor of 9

OpenStudy (anonymous):

how did you figure that out?

OpenStudy (anonymous):

d=14 ft r1=7 ft A1= pi*7^2 = 49*pi A2=49*3*pi = pi* r2^2 r2^2=147 => r2=12.12 ft circumference 2= 2*pi*r2 = 2*pi*12,12 = 24,24*pi circumference 1= 2*pi*r1= 2*pi*7= 14*pi 24.24/14= 1.731

OpenStudy (campbell_st):

oops no its a linear measurement the circumference will increase by a factor of root3

OpenStudy (anonymous):

thanks

OpenStudy (campbell_st):

sorry about the initial mistake...

OpenStudy (anonymous):

its ok